# Generative AI Beyond LLMs: System Implications of Multi-Modal Generation

### v1

@@ I Introduction
While image and video generation models have seen significant algorithm advancements in recent years, relatively little has been done to optimize the deployment of these models from a systems perspective. New system optimizations tailored towards system performance bottlenecks of TTI/TTV models have the potential to replace the generation of short video clips on the order of seconds with full movies and films. Other image generation tasks such as sticker-generation [9], educational material [10], and even scientific discoveries [11], could likewise benefit from system optimizations that enable increased speed and resolution. Overcoming systems challenges is critical to enabling future applications.
Figure 1: Across industry-scale datacenters, Text-to-Image (TTI) models use 14x more GPUs per model parameter during training and 1.4x higher memory utilization as compared to LLMs, demonstrating their growing importance at the datacenter scale.
To evaluate current state-of-the-art image/text generation, we first examine industry-scale generative deep learning tasks at the fleet-wide level. We find that while TTI/TTV models are an order of magnitude smaller than LLMs in terms of model parameters, the number of GPUs used for training is roughly in the same order-of-magnitude. In fact, the ratio of number of GPUs per model parameter is 14x higher for TTI models than LLMs, emphasizing the importance of running these models efficiently (Figure 1). Further fleet-wide characterization reveals that this emerging class of AI workloads has distinct system requirements — average memory utilization for TTI/TTV models is roughly 10% higher than LLMs.
@@
We additionally observe that traditional LLM paradigms such as Prefill/Decode and Sequence Length, do not map 1:1 onto TTI/TTV workloads. We profile sequence length over the course of inference and find that in contrast to LLMs, sequence length varies by up to 4x over the course of Diffusion model inference. In general, sequence length scales quadratically with image size in Diffusion models. Furthermore, we investigate the system performance bottleneck presented by Temporal Attention, which allows for cohesive frame generation across time in TTV models. This is in contrast to Spatial Attention, which attends across a 2D image. We find that Temporal Attention takes 2x the execution time as Spatial Attention, yet consumes 9x the FLOP count. We further observe the Temporal Attention bottleneck grows exponentially with number of frames, suggesting the need for future system optimizations.
Figure 2: Common Text-to-Image Model Architectures. Models consist of multiple independently-trained components, and are strung together during inference (shown here) to take text as input and generate an image output. Note that the top two models use a diffusion-based architectures (green), while bottom model uses a transformer-based architecture (red).
@@ II Understanding Multi-Modal Machine Learning Tasks
We first present a system-informed taxonomy of the current landscape of text-to-image (TTI) and text-to-video (TTV) generation models. Recent model development has seen a burification of architectures, each characterized by their own image generation process and system characteristics. Figure 2 illustrates the image generation pipeline for these two classes of workloads: (i) Diffusion Models, including both pixel- and latent- based models, and (ii) Transformer-based Models. These same two classes of workloads additionally form the fundamental building blocks of TTV models — since video generation models typically generate a series of images (i.e., frames) using a pretrained TTI model, and then ensure the frames are temporally consistent through additional temporal attention/convolutional layers. The subsequent discussion and analysis refers to the Inference pipeline of image/text generation. Unlike LLMs, TTI/TTV models consist of several different model components that are trained separately and then stiched together at inference time.
@@ II-A1 Diffusion Models
As shown in Figure 2, there are two distinct variations of diffusion models from a systems perspective — pixel and latent based. Pixel vs Latent models are distinguished by the parameter space of the diffusion model and the subsequent post-processing which is necessary once an initial image is generated. While pixel-based models operate the denoising process on standard image pixels, latent-based models transform the image into an embedding representation, making it more efficient for computation [17]. As a result, latent-based models can represent high-resolution images without the need to feed the image through subsequent SuperResolution (SR) networks as in the pixel case. This comes at the cost of a VAE or GAN-based deocder to convert latent space back to pixel space once finished.
@@ II-A2 Transformer Models
Figure 3: Detail on Diffusion and Transformer models. Note that Diffusion models consist of Resnet blocks, Self-Attention blocks, and Cross-Attention blocks while Transformer-based models contain Self-Attention, Cross-Attention, and FeedForward.
@@ II-B Text-to-Video Models
Figure 4: Pareto-Optimal curve showing tradeoff between model quality and system resources for various Text-to-Image models. Bottom left corner is optimal. Bolded points represent models further examined in model suite outlined in Section III.
@@ II-C System Design Space of Text-to-Image/Video Models
To further understand the system design space for emerging multi-modal generative AI technologies, Figure 4 illustrates state-of-the-art TTI generation technologies between the key
Model
Imagen [25]
Stable Diffusion[17]
Muse [21]
Parti [19]
Architecture
Diffusion (Pixel)
Diffusion (Latent)
Trans- former
Trans- former
Num Params
3B
1.45B
3B
20B
Num Layers
48
80
Model Dim
2048
4096
Attn Res
[32,16,8]
[4, 2, 1]
Text Cross
[32,16,8]
Attn Res
Channel Mult
[1,2,4,4]
[1,2,4,4]
Num Res Blocks
3
2
Per-Head Channels
64
8
Embed Dim
512
768
Compute
High
Medium
Low
Low
Memory
Medium
Low
Low
High
Latency
High
Medium
Low
Medium
TABLE I: Taxonomy of Text-to-Image Models
design dimensions of model accuracy (x-axis) and the number of trainable parameters (y-axis). Models are labeled as their citation number [18, 26, 27, 28, 29, 30, 21, 19, 31, 32, 33, 34, 17, 25, 35]. Accuracy is measured via previously reported values of FID score [36] on the COCO dataset [37]. We omit non-open source models since the lack of publicly available implementations prevents a more in-depth system analysis. As shown, the most optimal models tend towards the bottom left corner with lowest FID score and fewest number of parameters. In general, Diffusion models tend to have higher model quality for the same number of parameters as Transformer models. However, a diverse set of models lie on the Pareto-Optimal curve, including Imagen [25] (pixel-based), Stable Diffusion [17] (latent based) and Parti (transformer), the last of which offers higher model quality than diffusion models but at 4x the parameter count. Note that a corresponding figure for TTV models shows a similar trend. This exemplifies the importance of understanding system implications for these various architectures.
Figure 5: Text-to-Image/Video Models Roofline on A100 GPU. Diffusion models have higher arithmetic intensity than transformer-based TTI models, and fall in the compute-bound region. Transformer-based models tend to be memory-bandwidth bound at low batch sizes.
@@ III Experimental Methodology
Hardware Systems. We evaluate training and inference using NVIDIA A100 80GB GPUs. For inference analysis, we profile TTI/TTV model inference on a single GPU, since the model parameters can fit within the 80 GB memory constraints. When profiling model training, we use Fully Sharded Data Parallelism (FSDP) to train over multiple compute nodes, where each node consists of 8 A100 GPUs.
Tools. To perform our subsequent characterization and analysis, we use PyTorch Profiler to record execution timelines. We measure GPU kernel time and annotate the model to track which CPU operators correspond to a given GPU launch kernel. We develop a profiling framework to automate this process, via inserting hooks into the forward functions of the each module. We then develop scripts to efficiently parse the resulting Pytorch Profiler output and link GPU kernels to their corresponding annotation to determine operator breakdowns, speedup, etc. We construct a similar framework to analytically calculate FLOPs. We use the NVIDIA Nsight Compute tool to examine kernel breakdowns and analyze cache hit rates, memory utilization, etc.
@@ IV System Performance Characterization
TABLE 2: End-to-end speedup of Flash Attention as compared to Baseline Attention
@@ IV-A Analyzing System Performance Bottlenecks
The detailed operator time characterization reveals Attention is an important system performance bottleneck in baseline models — given that Attention consumes roughly 41.3% of execution time when averaged across the TTI/TTV model suite. To accelerate the execution time of the Amdahl’s law bottleneck, the recently proposed Flash Attention shows significant performance potential on GPUs [8]. The technique essentially allows for the large intermediate matrices produced in the Attention calculation to be tiled, thereby reducing memory accesses and offering significant speedup. Note that while Flash Attention was originally designed as a training optimization, previous work including the DeepSpeed Inference Framework and others has shown its benefit for inference as well [38]. Here we examine the impact of applying Flash Attention V2 across TTI/TTV inference workloads.
@@ IV-B Evaluating Effectiveness of Flash Attention Across TTI/TTV Model Suite
Training/ Prefill
1st token
Generate all pixels of the image at once
Process text prompt
Decode
2nd token
N/A
Generate each token autoregressively
@@
To understand the observed difference in Attention speedups across the model suite, we note that Attention kernel speedup varies as a factor of matrix size. We subsequently analyze TTI model speedup in the context of traditional LLM inference, which consists of two distinct phases: Prefill and Decode. Prefill (i.e., the initial processing of the prompt) allows for greater parallelization, as it is characterized by the processing of a large Nxd query vector, where N is sequence length and d is model dimension. A large NxN similarity matrix is created during the Attention calculation, which benefits from the tiling of Flash Attention due to reduced HBM memory accesses [8]. In contrast, the Decode stage generates tokens autoregressively, with queries of size 1xN. This generates a smaller Nx1 similarity matrix, which requires fewer memory accesses in the first place, and thus sees smaller speedup.
@@ V Impact of Sequence Length
Figure 7: Sequence length profiling across various models in model suite. Shown as sequence length over course of time. The variation in sequence length over time for diffusion models pose unique system constraints for these UNet-based models. Note sequence length of Stable Diffusion model actually goes up to 4096, but not shown here for plotting purposes.
@@ V-A Sequence Length Profiling
Figure 8: Frequency distribution of sequence lengths over the course of inference for Stable Diffusion model. Note the significance of the overlapping bars, where the value distribution shifts right for increasing image size and corresponds to the input/output size. The distribution associated with image size of 512 corresponds to Figure 7.
@@
We develop an analytical framework to model the changing memory and FLOPs requirements over the course of the forward pass of a Diffusion model.
We start with the desired image size $H_{O}$ x $W_{O}$, which is subsequently downsampled to latent space, $H_{L}$ x $W_{L}$. We define $text\_encode$ as the length of the encoded text prompt. Note that sequence length for the Self-Attention blocks in the UNet is based on the size of the latent-representation, and is governed by:
$(H_{L}*W_{L})\times(H_{L}*W_{L})$
while sequence length for the Cross-Attention blocks is additionally based on text encoding as shown:
$H_{L}\times W_{L}\times(text\_encode)$
@@
$2\times(H_{L}*W_{L})\times(H_{L}*W_{L})+2\times(H_{L}*W_{L})\times(text\_encode)$
@@
$\Biggr{[}2*\sum_{n=0}^{unetdepth-1}\frac{H_{L}W_{L}}{d^{n}}\Bigr{[}\frac{H_{L}W_{L}}{d^{n}}+text\_encode\Bigr{]}\Biggr{]}$
$+\frac{H_{L}W_{L}}{d^{unetdepth}}\times\Bigr{[}\frac{H_{L}W_{L}}{d^{unetdepth}}+text\_encode\Bigr{]}$
@@ V-B Implications of Scaling Image Size
We subsequently construct a case study on the Stable Diffusion model, in order to more concretely understand the impact of scaling image size (i.e., $H_{O}$, $W_{O}$). Figure 8 sweeps different input image sizes, and illustrates the corresponding sequence length distribution for Stable Diffusion inference. Note that as image size increases, sequence length distribution shifts to the right. By examining the case of a 512 x 512 image size, we note the distribution over sequence lengths is relatively equal, which confirms the symmetric nature of the sequence length profile for Stable Diffusion shown in Figure 7. As shown, the sequence lengths confine themselves to distinct buckets, which could allow future systems to tailor hardware towards sequence lengths of interest.
@@ VI-A Temporal versus Spatial Attention
We additionally find that Temporal Attention has a roughly 10x lower L1 cache hit rate for GEMM and softmax kernels as compared to Spatial Attention, as shown in Figure 12. This lower locality translates to a higher cache miss rate, thus increasing memory access latency time. In terms of L2 cache, GEMM kernels have roughly 10x lower hit rate. However, cache hit rates for elementwise and softmax remain the same or higher as the corresponding kernel in spatial attention.

### v2

@@ Abstract
Index Terms:
Generative AI, Multi-Modal, Diffusion Model, Transformer, Sequence Length, Attention
@@ I Introduction
While image and video generation models have seen significant algorithm advancements in recent years, relatively little has been done to optimize the deployment of these models from a systems perspective. New system optimizations tailored towards system performance bottleneck of TTI/TTV models have the potential to replace the generation of short video clips on the order of seconds with full movies and films. In this paper, we refer to the performance bottleneck as operators that dominate the execution time of the workload. Other image generation tasks such as sticker-generation [9], educational material [10], and even scientific discoveries [11], could likewise benefit from system optimizations that enable increased speed and resolution. Overcoming systems challenges is critical to enabling future applications.
Figure 1: Across industry-scale datacenters, Text-to-Image (TTI) models use roughly 14x more GPUs per model parameter during training and 1.35x higher memory utilization as compared to LLMs, demonstrating their growing importance at the datacenter scale.
To evaluate current state-of-the-art image/text generation, we first examine industry-scale generative deep learning tasks at the fleet-wide level. We find that while TTI/TTV models are an order of magnitude smaller than LLMs in terms of model parameters, the number of GPUs used for training is roughly in the same order-of-magnitude. In fact, the ratio of number of GPUs per model parameter is 14x higher for TTI models than LLMs, emphasizing the importance of running these models efficiently (Figure 1). Further fleet-wide characterization reveals that this emerging class of AI workloads has distinct system requirements — average memory utilization for TTI/TTV models is roughly 35% higher than LLMs.
@@
We additionally observe that traditional LLM paradigms such as Prefill/Decode and Sequence Length, do not map 1:1 onto TTI/TTV workloads. We profile sequence length over the course of inference and find that in contrast to LLMs, sequence length varies by up to 4x over the course of Diffusion model inference. In general, sequence length scales quadratically with image size in Diffusion models. Furthermore, we investigate the system performance bottleneck presented by Temporal Attention, which allows for cohesive frame generation across time in TTV models. This is in contrast to Spatial Attention, which attends across a 2D image. We find that Temporal Attention takes 2x the execution time as Spatial Attention, yet consumes 9x the FLOP count. We further observe the Temporal Attention bottleneck grows quadratically with number of frames, suggesting the need for future optimizations.
Figure 2: Common Text-to-Image Model Architectures. Models consist of multiple independently-trained components, and are strung together during inference (shown here) to take text as input and generate an image output. Note that the top two models use a diffusion-based architectures (green), while bottom two models use transformer-based architectures (red).
@@ II Understanding Multi-Modal Machine Learning Tasks
We first present a system-informed taxonomy of the current landscape of text-to-image (TTI) and text-to-video (TTV) generation models. Recent model development has seen a bifurcation of architectures, each characterized by their own image generation process and system characteristics. Figure 2 illustrates the image generation pipeline for these two classes of workloads: (i) Diffusion Models, including both pixel- and latent- based models, and (ii) Transformer-based Models. These same two classes of workloads additionally form the fundamental building blocks of TTV models — since video generation models typically generate a series of images (i.e., frames) using a pretrained TTI model, and then ensure the frames are temporally consistent through additional temporal attention/convolutional layers. The subsequent discussion and analysis refers to the Inference pipeline of image/text generation. Unlike LLMs, TTI/TTV models consist of several different model components that are trained separately and then stiched together at inference time.
@@ II-A1 Diffusion Models
As shown in Figure 2, there are two distinct variations of diffusion models from a systems perspective — pixel and latent based. Pixel vs Latent models are distinguished by the parameter space of the diffusion model and the subsequent post-processing which is necessary once an initial image is generated. While pixel-based models operate the denoising process on standard image pixels, latent-based models transform the image into an embedding representation, making it more efficient for computation [17]. As a result, latent-based models can represent high-resolution images without the need to feed the image through subsequent SuperResolution (SR) networks as in the pixel case. This comes at the cost of a VAE or GAN-based decoder to convert latent space back to pixel space once finished.
@@ II-A2 Transformer Models
Figure 3: Detail on Diffusion and Transformer models. Note that Diffusion models consist of Resnet blocks, Self-Attention blocks, and Cross-Attention blocks while Transformer-based models consider Self/Cross Attention and FeedForward.
@@ II-B Text-to-Video Models
Figure 4: Pareto-Optimal curve showing tradeoff between model quality and system resources for various Text-to-Image models. Bottom left corner is optimal. Bolded points represent models further examined in model suite outlined in Section III. Note that a corresponding figure for TTV models shows a similar trend. Diffusion TTV models often have both the lowest number of parameters and FID score.
Model Imagen [25] Stable Diffusion[17] Muse [21] Parti [19]
Architecture Diffusion (Pixel) Diffusion (Latent) Trans- former Trans- former
Num Params 3B 1.45B 3B 20B
Num Layers 48 80
Model Dim 2048 4096
Attn Res [32,16,8] [4, 2, 1]
Text Cross [32,16,8]
Attn Res
Channel Mult [1,2,4,4] [1,2,4,4]
Num Res Blocks 3 2
Per-Head Channels 64 8
Embed Dim 512 768
Compute High Medium Low Low
Memory Medium Low Low High
Latency High Medium Low Medium
TABLE I: Taxonomy of Text-to-Image Models
Figure 5: Text-to-Image/Video Models Roofline on A100 GPU. Diffusion models have higher arithmetic intensity than transformer-based TTI models, and fall in the compute-bound region. Transformer-based models are memory-bandwidth bound at low batch sizes.
@@ II-C System Design Space of Text-to-Image/Video Models
To further understand the system design space for emerging multi-modal generative AI technologies, Figure 4 illustrates state-of-the-art TTI generation technologies between the key design dimensions of model accuracy (x-axis) and the number of trainable parameters (y-axis). Models are labeled as their citation number [18, 26, 27, 28, 29, 30, 21, 19, 31, 32, 33, 34, 17, 25, 35]. Accuracy is measured via previously reported values of FID score [36] on the COCO dataset [37]. We omit non-open source models since the lack of publicly available implementations prevents a more in-depth system analysis. As shown, the most optimal models tend towards the bottom left corner with lowest FID score and fewest number of parameters. In general, Diffusion models tend to have higher model quality for the same number of parameters as Transformer models. However, a diverse set of models lie on the Pareto-Optimal curve, including Imagen [25] (pixel-based), Stable Diffusion [17] (latent based) and Parti (transformer), the last of which offers higher model quality than diffusion models but at 4x the parameter count. This exemplifies the importance of understanding system implications for these various architectures.
@@ III Experimental Methodology
Hardware Systems. We evaluate training and inference using NVIDIA A100 80GB GPUs. For inference analysis, we profile TTI/TTV model inference on a single GPU, since the model parameters can fit within the 80 GB memory constraints. When profiling model training, we use Fully Sharded Data Parallelism (FSDP) to train over multiple compute nodes, where each node consists of 8 A100 GPUs. Figure 6 shows how compute and communication scale as training is distributed over an increasing number of nodes.
Tools. To perform our subsequent characterization and analysis, we use PyTorch Profiler to record execution timelines. We measure GPU kernel time and annotate the model to track which CPU operators correspond to a given GPU launch kernel. We develop a profiling framework to automate this process, via inserting hooks into the forward pass. We then develop scripts to efficiently parse the resulting Pytorch Profiler output and link GPU kernels to their corresponding annotation to determine operator breakdowns, speedup, etc. We construct a similar framework to analytically calculate FLOPs. We use the NVIDIA Nsight Compute tool to examine kernel breakdowns and analyze cache hit rates, memory utilization, etc.
Figure 6: Compute versus communication scaling of Stable Diffusion model as training is distributed over an increasing number of nodes (i.e., 2, 8, and 16 A100 GPUs).
@@ IV System Performance Characterization
TABLE II: End-to-end speedup of Flash Attention as compared to Baseline Attention
@@ IV-A Analyzing System Performance Bottlenecks
The detailed operator time characterization reveals Attention is an important system performance bottleneck in baseline models — given that Attention consumes roughly 41.3% of execution time when averaged across the TTI/TTV model suite. To accelerate the execution time of Attention, the recently proposed Flash Attention shows significant performance potential on GPUs [8]. The technique essentially allows for the large intermediate matrices produced in the Attention calculation to be tiled, thereby reducing memory accesses and offering significant speedup. Note that while Flash Attention was originally designed as a training optimization, previous work including the DeepSpeed Inference Framework and others has shown its benefit for inference as well [38]. Here we examine the impact of applying Flash Attention V2 across TTI/TTV inference workloads.
@@ IV-B Evaluating Effectiveness of Flash Attention Across TTI/TTV Model Suite
Training/ Prefill 1st token Generate all pixels of the image at once Process text prompt
Decode 2nd token N/A Generate each token autoregressively
@@
To understand the observed difference in Attention speedups across the model suite, we note that Attention kernel speedup varies as a factor of matrix size. We subsequently analyze TTI model speedup in the context of traditional LLM inference, which consists of two distinct phases: Prefill and Decode. Prefill (i.e., the initial processing of the prompt) allows for greater parallelization, as it is characterized by the processing of a large Nxd query vector, where N is sequence length and d is model dimension. A large NxN similarity matrix is created during the Attention calculation, which benefits from the tiling of Flash Attention due to reduced HBM memory accesses [8]. In contrast, the Decode stage generates tokens autoregressively, with queries of size 1xN. This generates a smaller Nx1 similarity matrix, which requires fewer memory accesses in the first place, and thus sees smaller speedup. Broadly speaking, Prefill boils down to large matrix-matrix computations while Decode consists of matrix-vector computations.
@@ V-A Sequence Length Profiling
Figure 8: Sequence length profiling across various models in model suite. Shown as sequence length over course of time. The variation in sequence length over time for diffusion models pose unique system constraints for these UNet-based models. In particular, measured values of memory bandwidth utilization and last level cache (LLC) miss rate over the identical region of interest for Imagen and Parti models reveal they correlate with sequence length variation. Note sequence length of Stable Diffusion model actually goes up to 4096, but not shown here for plotting purposes.
Figure 9: Frequency distribution of sequence lengths over the course of inference for Stable Diffusion model. The value distribution shifts right for increasing image size, which corresponds to the input/output size. The distribution associated with image size of 512 corresponds to Figure 8.
@@
Note that we additionally quantify the memory bandwidth utilization and last level cache (LLC) miss rate for Imagen and Parti models over the same ROI as Figure 8. We find that memory bandwidth utilization and LLC miss rate follow the same pattern as the illustrated sequence length. This highlights the significance of sequence length variation over the course of model inference and its impact on the memory system.
We develop an analytical framework to model the changing memory and FLOPs requirements over the course of the forward pass of a Diffusion model. The model is built by statically examining the model computation to analyze the theoretical memory requirement. We then perform additional profiling runs in order to ensure accuracy of the model, and quantify matrix dimensions throughout model execution to verify the size of the matrices that our formulas predict.
We start with the desired image size $H_{O}\times W_{O}$, which is subsequently downsampled to latent space, $H_{L}\times W_{L}$. We define $text\_encode$ as the length of the encoded text prompt. Note that sequence length squared for the Self-Attention blocks in the UNet is based on the size of the latent-representation, and is governed by:
$(H_{L}\cdot W_{L})\cdot(H_{L}\cdot W_{L})$
while sequence length squared for the Cross-Attention blocks is additionally based on text encoding as shown:
$H_{L}\cdot W_{L}\cdot(text\_encode)$
@@
$2\cdot(H_{L}\cdot W_{L})\cdot(H_{L}\cdot W_{L})+2\cdot(H_{L}\cdot W_{L})\cdot(text\_encode)$
@@
$2\cdot\Biggr{[}2\sum_{n=0}^{unetdepth-1}\frac{H_{L}W_{L}}{d^{2n}}\Bigr{[}\frac{H_{L}W_{L}}{d^{2n}}+text\_encode\Bigr{]}$
$+\frac{H_{L}W_{L}}{d^{{2unetdepth}}}\Bigr{[}\frac{H_{L}W_{L}}{d^{{2unetdepth}}}+text\_encode\Bigr{]}\Biggr{]}$
@@ V-B Implications of Scaling Image Size
We subsequently construct a case study on the Stable Diffusion model, in order to more concretely understand the impact of scaling image size (i.e., $H\_O$, $W\_O$). Figure 9 sweeps different input image sizes, and illustrates the corresponding sequence length distribution for Stable Diffusion inference. Note that as image size increases, sequence length distribution shifts to the right. By examining the case of a 512 x 512 image size, we note the distribution over sequence lengths is relatively equal, which confirms the symmetric nature of the sequence length profile for Stable Diffusion shown in Figure 8. As shown, the sequence lengths confine themselves to distinct buckets, which could allow future systems to tailor hardware towards sequence lengths of interest.
@@ VI-A Temporal versus Spatial Attention
We profile a variety of performance counters to further compare Temporal and Spatial Attention, including Misses Per Thousand Instructions (MPKI), Average Memory Access Time (AMAT), and Memory Bandwidth. As Figure 13 shows, we find that GEMM and elementwise kernels have a 1.96x and 1.22x higher L2 MPKI for Temporal attention as compared to Spatial Attention. We additionally find the AMAT for GEMM, elementwise, and softmax kernels is 28, 4, and 18 cycles longer for Temporal Attention, respectively. By collecting an ensemble of performance counters, we make progress towards understanding this system bottleneck, yet there is more work to be done to fully understand the nature of Temporal Attention.